OpenStudy (anonymous):

Use a graphing utility to sketch the curve represented by the parametric equations (indicate the direction of the curve). Then eliminate the parameter and write the corresponding rectangular equation whose graph represents the curve. x = 4 + 2 cos theta y= -1 + 2 sin theta

OpenStudy (asnaseer):

you can use wolframalpha to graph this, see here: http://www.wolframalpha.com/input/?i=x+%3D+4+%2B+2*cos%28theta%29%2C+y%3D+-1+%2B+2*sin%28theta%29

OpenStudy (asnaseer):

to convert to rectangular coordinates, just rearrange the expressions so you get cos(theta) equals something and sin(theta) equals something, then square each one and add - that will eliminate te terms in theta since:\[\cos^2(\theta)+\sin^2(\theta)=1\]

OpenStudy (anonymous):

Thank you <3 ^^

OpenStudy (asnaseer):

yw :)